Rewrite cos 4x + cos 2x as a product and simplify if possible.  
Answer #1

= = Now, cosyre e cobar 2cos 82x -1+ Cosak E: 68.436 = 26822x-] 260822x + 2008an Cosar -1 = 200320 (Coszx+) -(682x41) = (COB2